OpenStudy (anonymous):

Solve: lim x->infinity (x/(1+x))^x

OpenStudy (anonymous):

To Clarify: \[\lim_{x \rightarrow \infty}\left( \frac{ x }{ x +1} \right)^{x} \]

OpenStudy (anonymous):

which could go to \[\lim_{x \rightarrow \infty}\epsilon^{xlnx - \ln(x+1) }\]

OpenStudy (anonymous):

\[\lim_{x\to\infty}\left(\frac{x}{x+1}\right)^x=\lim_{x\to\infty}\exp\left[\ln\left[\left(\frac{x}{x+1}\right)^x\right]\right]\] Using the continuity of the exponential function, you have \[\exp\left[\lim_{x\to\infty}x\ln\left(\frac{x}{x+1}\right)\right]\] The limit above gives you the indeterminate form \(\infty\cdot0\), so rewrite a bit to get \[\exp\left[\lim_{x\to\infty}\frac{\ln\left(\frac{x}{x+1}\right)}{\frac{1}{x}}\right]\] Now, the limit gives you indeterminate form \(\dfrac{0}{0}\). Apply L'Hopital's rule as many times as necessary.

OpenStudy (anonymous):

Ok, I know the answer is \[e ^{-1}\].. Will give it a shot. Thanks man.

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

L'hopitals, take the limit of the same function. Just denominator and numerator derived yes?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\[\lim_{x\to\infty}\frac{f(x)}{g(x)}=\lim_{x\to\infty}\frac{f'(x)}{g'(x)}\] given that the left limit is 0/0 or ±∞/∞.

OpenStudy (anonymous):

so this would give me exp \[\exp\left[ \lim_{x \rightarrow \infty} \frac{ x }{ 1+x }\right]\]

OpenStudy (anonymous):

Which would be e^0 which is wrong..

OpenStudy (zarkon):

it would be \(e^1\) which would still be wrong. ;)

OpenStudy (anonymous):

My second mistake then, Yay :P

OpenStudy (anonymous):

Numerator:\[\frac{d}{dx}\ln\left(\frac{x}{x+1}\right)=\frac{\frac{x+1-x}{(x+1)^2}}{\frac{x}{x+1}}=\frac{1}{x(x+1)}\] Denominator: \[\frac{d}{dx}\frac{1}{x}=-\frac{1}{x^2}\] \[\frac{\frac{1}{x(x+1)}}{-\frac{1}{x^2}}=\color{red}-\frac{x}{x+1}\]

OpenStudy (anonymous):

So thats where the neg is hiding, thank you :)

OpenStudy (anonymous):

One last question, \[\lim_{x \rightarrow \infty } -\frac{ x }{ x + 1}\] is -1 because the +1 in the denominator becomes negligible yes?

OpenStudy (anonymous):

You could think of it that way, yeah. If you divide by each term in numerator and denominator by \(x\), you have \[-\frac{1}{1+\frac{1}{x}}\] The \(\dfrac{1}{x}\) in the denominator approaches 0 as x approaches infinity.